The following is a 90% confidence interval for p:(0.26, 0.54). How large was the sample used to construct this interval?
Answer #1

Solution :

given that

Lower confidence interval = 0.26

Upper confidence interval = 0.54

Point estimate = \hat p = (Lower confidence interval + Upper confidence interval ) / 2

=  \hat p = (0.26+0.54) / 2

Point estimate = \hat p =0.4

Margin of error = E = Upper confidence interval - \hat p

=0.54 - 0.4

Margin of error = E = 0.14

Solution :

Given that,

\hat p = 0.4

1 - \hat p = 1 - 0.4 = 0.6

margin of error = E = 0.14

At 90% confidence level z

\alpha = 1 - 90%  

\alpha = 1 - 0.90 =0.10

\alpha/2 = 0.05

Z\alpha/2 = Z0.05 = 1.645

Sample size = n = (Z\alpha/2 / E)2 * \hat p * (1 - \hat p )

= (1.645 / 0.14)2 * 0.4 * 0.6

=33